OpenStudy (anonymous):

Solve K = (b – 2a) for b. You must show all work.

OpenStudy (anonymous):

we need to get b by itself on one side so first is to add 2a to both sides: K = b - 2a K + 2a = b - 2a + 2a K + 2a = b that wasn't too bad eh?

OpenStudy (anonymous):

thanks i was confused with all the letters thks:)

OpenStudy (anonymous):

yw...

OpenStudy (anonymous):

Just remember that the letters are just filling in for numbers you will put in later ;)